# Principal stratification with recurrent events truncated by a terminal event: A nested Bayesian nonparametric approach

### v2

@@ Abstract
Recurrent events often serve as key endpoints in clinical studies but may be prematurely truncated by terminal events such as death, creating selection bias and complicating causal inference. To address this challenge, we propose novel causal estimands within the principal stratification framework, introducing a refined always-survivor stratum that defines survival until the final recurrent event rather than a fixed time point, yielding more stable causal contrasts. We develop a flexible Bayesian nonparametric prior—the enriched dependent Dirichlet process—specifically designed for joint modeling of recurrent and terminal events, addressing a limitation where standard Dirichlet process priors create random partitions dominated by recurrent events, yielding poor predictive performance for terminal events. Our nested structure separates within-arm and cross-arm dependence through a dual-frailty framework, enabling transparent sensitivity analysis for non-identifiable parameters. Simulations are carried out to show that our method has superior performance compared to existing methods. We also illustrate the proposed Bayesian methods to infer the causal effect of intensive blood pressure control on recurrent cardiovascular events in a cardiovascular clinical trial.
Keywords: Causal inference; Bayesian nonparametric methods; Principal stratification; Semi-competing risks; Enriched dependent Dirichlet process
@@ 1 Introduction
Recurrent events are commonly observed in clinical trials and observational studies. Examples include hospitalizations or exacerbations of illness, and recurrent endpoints can provide valuable insights into the disease burden. Thus, the ability to draw valid causal inference about treatment effects on recurrent events can facilitate informed decision-making. Nonetheless, analyzing recurrent events becomes complex in the presence of a terminal survival event, since the death can introduce selection bias, because patients who die under one treatment condition may systematically differ in their baseline risk profiles from those who remain alive. In such cases, standard regression analyses may yield inaccurate conclusions about the treatment efficacy, since the estimated effects may primarily reflect differences in susceptible risk profiles rather than genuine causal effects on the recurrent process.
@@
Principal stratification (Frangakis2002) is a general framework for causal inference that involves post-treatment intermediate variables and has gained prominence in addressing complexities caused by truncation due to death. Within this framework, subpopulations—known as principal strata—are defined by intermediate potential outcomes under each treatment condition, enabling researchers to define unambiguous causal effects within these subgroups. For example, the always-survivor subpopulation, consisting of patients who would survive until the non-terminal event occurs regardless of treatment assignment, was first introduced by Robins1986 and formally established by Zhang_Rubin_2003. The associated identifiability assumptions have been previously studied when a non-mortality outcome is truncated by death (e.g., Zhang_Rubin_2003).
@@
Leveraging the principal stratification framework, a flourishing line of research has focused on evaluating the causal effect of treatment on a non-recurrent time-to-event outcome subject to semi-competing risks by death (Comment2019; Xu2022; Nevo2022). By contrast, much fewer attention was devoted to identifying causal effects on recurrent events subject to semi-competing risks by death, in which the recurrent event process may be informatively truncated by death. An exception is Lyu2023, who proposed a causal estimand for the average number of recurrent events among always-survivors and developed a Bayesian parametric joint modeling approach to simultaneously represent the recurrent event and death processes. In their formulation, a single time index $t$ was used to define both the number of recurrent events and the always-survivor stratum. However, adopting a single time point $t$ to define causal estimands for recurrent event analysis with a terminal event can be less ideal because it considers the always-survivor population only at that specific time. This perspective makes the subpopulation under study vary by $t$, overlooking how treatment effects may accumulate or change over time for a fixed subpopulation and impeding comparisons of estimands across different time points. Additionally, this traditional definition of always-survivors may unnecessarily exclude units from the always-survivor stratum in the context of recurrent events. Finally, principal stratification typically requires careful parametric model specification that encompasses stringent regression assumptions and prior distributions. Therefore, the parametric modeling approach (e.g., Comment2019; Lyu2023) could be susceptible to model misspecification bias, especially under potentially complex data generating processes.
@@
To address these challenges, we develop a nested Bayesian nonparametric approach for causal inference with recurrent events in the presence of a terminal event. Our contributions to the literature are several-folded. First, we propose new causal estimands tailored for recurrent event analyses subject to semi-competing risks by death. We explain the limitations of the existing definition of always-survivors for causal recurrent event analysis and introduce refined estimands that are more robust to the choice of time index and better capture treatment effects over time. We then consider a sensitivity approach to derive the identification results under a principal stratification framework. In addition to identification, our second contribution is to develop a flexible Bayesian nonparametric (BNP) approach for inference. BNP approaches have been increasingly used for recurrent event analysis to mitigate biases arising from model misspecification, a concern in many existing parametric approaches. In both causal and non-causal contexts, previous efforts have employed Dirichlet process (DP) priors to jointly model the recurrent event and terminal event processes (e.g., Paulon2020; Tian2024). We discuss in detail a potential pitfall wherein a commonly used DP prior induces a random partition primarily determined by the recurrent events, underestimating the effect of the terminal event. As an improvement, we introduce a nested structure in the DP prior and incorporate dependence on covariates to relax the stringent exchangeability assumption. We refer to our new prior as the enriched dependent Dirichlet process (EDDP) prior, and we develop a fully tractable, efficient Gibbs sampling algorithm for implementation. Third, we introduce a joint‑frailty framework for causal inference that cleanly separates within‑arm (identifiable) and cross‑arm (non-identifiable) dependence. This is in contrast to existing frailty-based causal methods (e.g., Comment2019; Lyu2023), which assume a single common frailty, thereby conflating two distinct layers of dependence: (i) the within-arm dependence between recurrent and survival events, and (ii) the cross-arm dependence between potential outcomes under treatment and control. In our approach, each individual instead carries two latent frailties, one per each treatment arm, whose marginal distributions are estimated nonparametrically, while their unidentifiable correlation is treated as an interpretable sensitivity parameter. By varying the sensitivity parameter within a Bayesian g-computation scheme, we transparently track how causal conclusions might shift under different cross‑world dependence structures. Finally, through simulations, we demonstrate that our methods outperform existing approaches in terms of accuracy and robustness. Our proposed methodology is illustrated by re-analyzing recurrent event data from a randomized clinical trial.
1.1 A motivating example
The Systolic Blood Pressure Intervention Trial (SPRINT) is a randomized clinical trial that tested whether targeting a systolic blood pressure (SBP) below $120$ mm Hg, rather than the conventional goal of below $140$ mm Hg, would reduce cardiovascular morbidity and mortality in adults at elevated cardiovascular risk (SPRINT_2015). A total of $9361$ participants aged $50$ years or older were randomly assigned in a 1:1 ratio to an intensive-treatment strategy—monthly medication adjustments to achieve SBP <120<120 mm Hg—or a standard-treatment strategy involving quarterly visits to maintain SBP 135135139139 mm Hg under a prespecified antihypertensive formulary.
In SPRINT, every occurrence of the cardiovascular events that define the primary composite outcome is date-stamped. These events are myocardial infarction, non-MI acute coronary syndrome, stroke, heart-failure decompensation requiring intravenous therapy in a hospital or emergency setting, and cardiovascular death attributed to coronary, cerebrovascular, or other vascular causes. Survival endpoints further include the composite cardiovascular outcome and all-cause mortality. For causal analysis of the SPRINT trial, we use both the counts and timing of non-fatal cardiovascular events to measure the burden of adverse encounters under each SBP target regimen. By tracking each participant’s sequence of events alongside the terminal event, we are interested in the causal comparison between the intensive and standard treatment arms on the total number of events as well as the rate of event occurrence. To mitigate bias introduced by death, we concentrate on the treatment effect within subpopulations whose recurrent events are not truncated by the terminal event.
@@ 2 Notation, Setup, and Estimands
We consider a study involving $n$ units, each assigned to one of two treatment arms at the beginning of the study: $Z_{i}=1$ for the treatment group and $Z_{i}=0$ for the control group. For unit $i$, let $D_{i}^{z}$ denote the potential time to death under treatment $z\in\{0,1\}$. The observed survival time is denoted by $\mathcal{T}_{i}=\min(D_{i}^{Z_{i}},C_{i}^{Z_{i}})$, where $C_{i}^{Z_{i}}$ denotes the potential censoring time due to loss to follow-up for any reason other than the survival event under the assigned treatment. The event indicators $\delta_{i}^{D}=\mathbbm{1}\{\mathcal{T}_{i}=D_{i}^{Z_{i}}\}$ and $\delta_{i}^{C}=\mathbbm{1}\{\mathcal{T}_{i}=C_{i}^{Z_{i}}\}$ represent that the observed survival time corresponds to death ($\delta_{i}^{D}=1$) or loss to follow up censoring ($\delta_{i}^{C}=1$), respectively. We assume that the potential cumulative number of recurrent events for subject $i$ under treatment $z$, $N_{i}^{z}(\cdot)$, follows a point process. The potential point process $N_{i}^{z}(t)$ is well-defined only for $0\leq t<D_{i}^{z}$ and undefined for $t\geq D_{i}^{z}$ due to truncation (that is, we use $N_{i}^{z}(t)=*$ for the undefined case). Let $N_{i}=N_{i}^{Z_{i}}(\mathcal{T}_{i})$ denote the observed number of recurrent events until the last follow-up. Let $T_{ij}^{z}$ represent the time of the $j$-th recurrent event for subject $i$ under treatment $z$ for any $j\in\mathbb{N}$. Since the recurrent event is truncated for those who experince the terminal event, we define $T_{ij}^{z}=*$ for all $j$ with $T_{ij}^{z}>D_{i}^{z}$. As a special case, we write $T^{z}_{i0}=0$. The observed event times are $T_{ij}=T_{ij}^{Z_{i}}$, which are observed only for $j$ with $T_{ij}\leq\mathcal{T}_{i}$. The observed event times are $\{T_{ij}\}_{j=1}^{N_{i}}$, and the gap time between two successive events is defined as $W_{i1}^{z}=T_{i1}^{z}$ and $W_{ij}^{z}=T_{ij}^{z}-T_{i(j-1)}^{z},$ for $j\geq 2$. For the observed data, the gap times are $W_{ij}=T_{ij}-T_{i(j-1)}$. Finally, let $\mathbf{X}_{i}\in\mathcal{X}$ denote a vector of baseline covariates for unit $i$. The observed data for each unit then consist of the tuple $O_{i}=\{\mathcal{T}_{i},\delta_{i}^{D},\delta_{i}^{C},N_{i},\{T_{ij}\}_{j=1}^{N_{i}},Z_{i},\mathbf{X}_{i}\}$. We consider the following identification assumptions.
Assumption 1 (Consistency).
$N_{i}(t)=Z_{i}N_{i}^{1}(t)+(1-Z_{i})N_{i}^{0}(t)$.
@@
$Z_{i}\perp\{N_{i}^{0}(t),D_{i}^{0},N_{i}^{1}(t),D_{i}^{1}\}\mid\mathbf{X}_{i}.$
@@
Assumptions 1 and 2 are standard in randomized trials and observational studies. The first condition assumes away interference and the second condition assumes away any unmeasured confounders, and is guaranteed to hold under randomization. Assumption 3 states that potential censoring times are conditionally independent of the potential event outcomes, and is a common assumption in causal survival analysis. Although empirically unverifiable due to the cross-world counterfactuals, this assumption is reasonable in SPRINT because the majority of right-censored events were due to pre-specified end of follow-up.
2.1 Principal stratification in continuous time

### v3

@@ Abstract
Recurrent events often serve as key endpoints in clinical studies but may be prematurely truncated by terminal events such as death, creating selection bias and complicating causal inference. To address this challenge, we develop a Bayesian nonparametric framework to address potential selection bias due to truncation by death within the continuous-time principal stratification framework. We introduce causal estimands for recurrent events in the presence of a terminal event and derive a partial identification result for the estimand under a dual-frailty framework, enabling transparent sensitivity analysis for non-identifiable parameters. We then propose a flexible Bayesian nonparametric prior, the enriched dependent Dirichlet process, specifically designed for joint modeling of recurrent and terminal events, addressing a limitation where standard Dirichlet process priors create random partitions dominated by recurrent events, yielding poor predictive performance for terminal events. Simulations are carried out to show that our method has superior performance compared to existing methods. We apply the proposed new Bayesian nonparametric methods to infer the causal effect of a structured exercise program on rehospitalizations, which are subject to truncation by death.
Keywords: Chronic Heart Failure; Causal inference; Bayesian nonparametric methods; Principal stratification; Recurrent event analysis; Enriched dependent Dirichlet process
@@ 1 Introduction
Clinical management of chronic heart failure often extends beyond guideline-directed medical therapy to include nonpharmacologic strategies such as exercise training and lifestyle modification. For stable patients with chronic systolic heart failure, exercise training is frequently recommended, but clinicians must weigh the feasibility, tolerance, and burden of a structured exercise prescription relative to potential benefits on clinically meaningful outcomes, in particular heart-failure hospitalizations and survival (Heidenreich et al., 2022). The HF-ACTION (Heart Failure: A Controlled Trial Investigating Outcomes of Exercise Training) trial was designed to address this question by testing whether adding supervised aerobic exercise training to usual care reduces all-cause mortality and hospitalization in patients with chronic systolic heart failure (O’Connor et al., 2009). In practice, HF-ACTION speaks to a common outpatient dilemma: whether a formal exercise program can deliver measurable benefit on morbidity and mortality, without unacceptable safety concerns or poor adherence that limits effectiveness.
Addressing this question requires moving beyond time-to-first hospitalization to quantifying how exercise affects the overall burden and timing of repeat hospital admissions. For chronic heart failure, patients commonly experience multiple rehospitalizations, so analyses restricted to a single “first event” can miss clinically meaningful differences in cumulative morbidity and disease burden. The reduction in rehospitalization burden is also economically meaningful, because heart-failure hospitalizations are a major driver of overall spending (Reed et al., 2010) and even modest decreases in recurrent admissions could translate into substantial per-patient and population-level cost savings. However, assessing the cumulative burden of rehospitalization can be complicated when the intervention also influences survival: longer survival increases exposure time for future hospitalizations, while earlier death both reduces time at risk and truncates the observation period for recurrent events like rehospitalization. Further, when mortality differs by treatment, the observed recurrent-event process is observed under differential truncation, and comparisons can be distorted because the set of individuals who remain under observation later in follow-up is treatment-dependent. As a result, conventional regression approaches may conflate a treatment effect on the recurrent-event process with shifts in the composition of survivors, yielding estimates that reflect differential selection rather than a causal effect on the cumulative burden of rehospitalization. HF-ACTION exemplifies this broader challenge in evaluating chronic-disease management, where observed rehospitalizations are survivor-dependent while subject to differential truncation.
While capturing the full history of hospitalizations provides a more complete picture of disease burden, analyzing such recurrent event processes introduces significant methodological challenges. The central issue here is how to define the causal effect of interest. In the presence of death, the phrase “effect on rehospitalization” is inherently ambiguous unless the target population and the counterfactual comparisons are clearly specified. A naive comparison of observed rehospitalization counts or rates between treatment arms fails to isolate the effect of exercise on the recurrent-event process, primarily because the treatment may simultaneously alter survival and, by extension, the opportunity to experience subsequent hospitalizations. Consequently, such comparisons conflate distinct phenomena: differences in survival time, differences in the frequency or timing of rehospitalization while alive, and shifts in the composition of the surviving cohort over time. To obtain a clinically interpretable causal effect, it is therefore necessary to explicitly define a common subpopulation for whom the rehospitalization burden is being compared and the specific counterfactual scenarios under which such a comparison remains meaningful.
@@
Principal stratification (Frangakis and Rubin, 2002) provides a general framework for causal inference that involves post-treatment intermediate variables and has gained prominence in addressing complexities caused by truncation due to death. Within this framework, subpopulations, known as principal strata, are defined by intermediate potential outcomes under each treatment condition, enabling researchers to define unambiguous causal effects within these subgroups. For example, the always-survivor subpopulation, consisting of patients who would survive until the non-terminal event occurs regardless of treatment assignment, was first introduced by Robins (1986) and formally established by Zhang and Rubin (2003). The associated identifiability has been previously studied when a non-mortality outcome is truncated by death, where death was typically considered as a simple, dichotomous event (e.g., Long and Hudgens, 2013; Tchetgen Tchetgen, 2014).
@@
Extending the principal stratification framework to continuous time, a flourishing line of research has focused on evaluating the causal effect of treatment on a non-recurrent time-to-event outcome subject to semi-competing risks by time to death, or more generally, time to a terminal event (Comment et al., 2025; Xu et al., 2022; Nevo and Gorfine, 2022). By contrast, much less attention has been devoted to identifying causal effects on recurrent events subject to semi-competing risks by death, in which the recurrent event process may be informatively truncated by death. An exception is Lyu et al. (2023), who proposed a causal estimand for the average number of recurrent events among always-survivors and developed a Bayesian parametric joint modeling approach to simultaneously represent the recurrent event and death processes. In their formulation, a single time index $t$ was used to define both the number of recurrent events and the always-survivor stratum. However, adopting a single time point $t$ to define causal estimands for recurrent event analysis with a terminal event can be less ideal because it considers the always-survivor population only at that specific time. This perspective makes the subpopulation under study vary by $t$, overlooking how treatment effects may accumulate or change over time for a fixed subpopulation and impeding comparisons of estimands across different time points. Finally, principal stratification typically requires careful parametric model specification that encompasses stringent regression assumptions and prior distributions. Therefore, the parametric modeling approach (e.g., Nevo and Gorfine, 2022; Lyu et al., 2023; Comment et al., 2025) could be easily susceptible to model misspecification bias, especially under potentially complex data-generating processes.
@@
To address these challenges, we develop a Bayesian nonparametric framework for causal inference with recurrent events in the presence of a terminal event. Our novel contributions are severalfold. First, we propose new causal estimands and their identification results tailored for recurrent event analyses under semi-competing risks with death. Building on the double-index approach of Comment et al. (2025) to define always-survivor stratum, we derive partial identification results under a dual-frailty framework that cleanly separates within-arm (identifiable) dependence from cross-arm (non-identifiable) dependence. In particular, cross-world dependence is indexed by an interpretable sensitivity parameter, enabling a transparent sensitivity analysis of the cross-world dependence structure. Second, we develop a flexible Bayesian nonparametric (BNP) approach for inference. BNP methods have been increasingly used in recurrent event analysis to mitigate biases from model misspecification, a concern in many parametric approaches. In both causal and non-causal settings, previous work has employed Dirichlet process (DP) priors to jointly model recurrent and terminal event processes (e.g., Paulon et al., 2020; Tian et al., 2024). We discuss a potential pitfall in which a commonly used DP prior induces a random partition driven primarily by recurrent events, leading to inadequate learning about the terminal event component. To address this limitation, we introduce a nested DP structure and incorporate covariate dependence to relax the stringent exchangeability assumption. We refer to the resulting prior as the enriched dependent Dirichlet process (EDDP), which also naturally accommodates the dual-frailty framework. We also develop a fully tractable and efficient Gibbs sampling algorithm for posterior computation and demonstrate through simulation studies that our method outperforms existing approaches in accuracy and robustness. Finally, we analyze the HF-ACTION trial using the proposed methodology. Our analysis suggests that the formal exercise training reduces the expected number of rehospitalizations over long-term follow-up, with treatment effects becoming more pronounced over time. We show that the proposed estimand captures clinically relevant features of treatment benefit that conventional causal estimands do not fully reflect in the HF-ACTION setting. In addition, by varying the sensitivity parameter within a Bayesian g-computation scheme, we transparently assess how causal conclusions change under alternative cross-world dependence structures. The main conclusions remain stable across small to moderately large values of the sensitivity parameter. Only under an extreme dependence assumption, where the proposed dual-frailty formulation effectively collapses to a single-frailty approach, does the evidence become attenuated, highlighting the importance of modeling distinct cross-world frailties in this setting.
1.1 HF-ACTION (Heart Failure: A Controlled Trial Investigating Outcomes of Exercise Training)
HF-ACTION was a multicenter randomized clinical trial designed to evaluate whether adding a structured aerobic exercise program to guideline-based medical therapy reduces clinical events in patients with chronic systolic heart failure (O’Connor et al., 2009). Medically stable outpatients with left ventricular ejection fraction (LVEF) $\leq 35\%$ and New York Heart Association (NYHA) class II–IV symptoms despite optimal therapy were randomized 1:1 to usual care alone or usual care plus guided aerobic exercise training. The exercise intervention consisted of a structured supervised phase, with a goal of 36 sessions over approximately 3 months, followed by continued home-based training with ongoing adherence monitoring and encouragement during follow-up; the protocol does not indicate that exercise training stopped after a first rehospitalization. The usual-care arm did not receive a formal exercise prescription. In the original HF-ACTION trial, the primary endpoint was the composite of all-cause mortality or all-cause hospitalization, analyzed as a time-to-first-event outcome, with prespecified secondary endpoints of all-cause mortality, cardiovascular mortality or cardiovascular hospitalization, and cardiovascular mortality or heart-failure hospitalization. Thus, our focus on rehospitalization burden extends the original trial perspective beyond first-event analysis. In this work, the outcome of interest is the sequence of hospitalization events recorded during 4 years of follow-up, and death is treated as an absorbing terminal event that truncates observation of subsequent hospitalizations. This structure motivates estimands that go beyond time-to-first-event: for chronic heart failure, the cumulative burden and timing of repeated hospitalizations are clinically meaningful, but are entangled with survival because longer survival both increases time at risk for hospitalization and reduces truncation of the recurrent process by death.
Table 1 summarizes baseline covariates, events during follow-up, and distribution of rehospitalization by treatment arms in the analytic dataset. As expected in a randomized trial, baseline characteristics are well balanced across arms, including key prognostic factors such as left ventricular ejection fraction (LVEF), New York Heart Association (NYHA) class, and cardiopulmonary exercise (CPX) test duration. Mean follow-up is 928 days averaged across both arms, reflecting common trial protocol and administrative censoring. Hospitalizations are common and recurrent: more than 50% of participants experience at least one hospitalization, and the distribution of rehospitalization exhibits both excess zeros and a right tail (maximum 12 events). The gap-time summaries in Table 1 show substantial heterogeneity in the spacing of subsequent events, but the gap times decrease sharply for later recurrences. As a result, inference about late-event dynamics will be driven by a small subset with frequent hospitalizations.
Table 2 focuses on the clinically important subset of participants with at least one hospitalization. Two patterns are especially relevant for our analyses. First, within this subset, death occurs in a substantial portion of participants: more than 20% die during follow-up, suggesting that the rehospitalization process may beinformatively truncated by mortality among higher-risk patients. Second, among decedents the last observed hospitalization often occurs closer to death than the corresponding time from the last hospitalization to non-death censoring among survivors, reinforcing the interdependence of rehospitalization burden and mortality.
These empirical features highlight why causal analysis of rehospitalizations in HF-ACTION cannot rely solely on naive comparisons of observed hospitalization counts or rates between treatment arms, even under randomization. Because exercise training may affect both hospitalization and survival, and death censors future hospitalizations, the observed recurrent-event process mixes (i) the causal effect of treatment on hospitalization dynamics while alive and (ii) the causal effect of treatment on survival (and hence on time at risk and truncation). To separate these mechanisms, our methodology adopts a principal stratification perspective tailored to recurrent outcomes: we target treatment effects within a subpopulation whose recurrent histories up to a chosen horizon $r$ would not be truncated by death under either treatment assignment. For inference, we couple this estimand with a flexible Bayesian framework that jointly models hospitalization gap times and the terminal event while accommodating strong heterogeneity in event processes. Finally, because cross-arm dependence between potential outcomes is fundamentally unidentifiable, we employ a joint-frailty sensitivity formulation that separates identifiable within-arm dependence (between hospitalization and death) from non-identifiable cross-arm dependence, enabling transparent assessment of how causal conclusions shift under alternative cross-world dependence structures.
Table 1: Baseline covariates, follow-up, and hospitalization (recurrent-event) summaries by treatment arm in the HF-ACTION trial. Binary covariates and count-based summaries are shown as $n$ (%), continuous summaries as mean (SD) or median [Q1,Q3]. Hospitalization and follow-up times are measured in days from randomization. Gap-time summaries condition on having at least $j$ events and report the gap ending in the $j$th event.
Characteristic Usual care (n=1070) Exercise training (n=1060) Baseline covariates Age (years) 58.51 (12.97) 58.60 (12.38) Female 281 (26.3%) 318 (30.0%) BMI (kg/m2) 31.04 (7.08) 30.95 (7.16) NYHA class III/IV 379 (35.4%) 400 (37.7%) LVEF (%) 25.23 (7.36) 25.19 (7.60) Ischemic etiology 546 (51.0%) 550 (51.9%) Angina 250 (23.4%) 277 (26.1%) Prior MI 455 (42.5%) 444 (41.9%) Stroke history 113 (10.6%) 105 (9.9%) History of diabetes 335 (31.3%) 343 (32.4%) History of depression 237 (22.1%) 215 (20.3%) Systolic BP (mm Hg) 113.98 (18.57) 113.57 (17.92) Diastolic BP (mm Hg) 70.15 (11.26) 70.27 (11.32) Baseline heart rate (bpm) 70.82 (11.66) 70.95 (11.25) CPX test duration 9.88 (4.03) 9.74 (3.80) Follow-up and event summaries Follow-up time (days) 923.9 (400.8) 931.3 (388.3) Death observed 183 (17.1%) 168 (15.8%) Any hospitalization 594 (55.5%) 561 (52.9%) Hospitalizations per participant 0.922 (1.198) 0.868 (1.108) Total hospitalizations 987 920 Hospitalization rate (per 100 person-years) 36.46 34.04 Days to first hospitalization among those with $\geq 1$ event 293 [128, 573] 275 [120, 544] Distribution of recurrent events per participant 0 hospitalizations 476 (44.49%) 499 (47.08%) 1 hospitalization 360 (33.64%) 341 (32.17%) 2 hospitalizations 143 (13.36%) 132 (12.45%) 3 hospitalizations 64 (5.98%) 62 (5.85%) 4 hospitalizations 12 (1.12%) 11 (1.04%) 5 hospitalizations 7 (0.65%) 9 (0.85%) 6 hospitalizations 1 (0.09%) 3 (0.28%) 7+ hospitalizations 7 (0.65%) 3 (0.28%) Gap times between recurrent events (days) Gap between 1st and 2nd events 132 [39, 333] 107 [32, 363] Gap between 2nd and 3rd events 138 [58, 292] 106 [35, 244] Gap between 3rd and 4th events 127 [49, 195] 86 [49, 128] Gap between 4th and 5th events 99 [64, 284] 67 [46, 200] Gap to 6th+ event (from previous event) 78 [38, 100] 60 [26, 94]
Table 2: Summary among participants with at least one observed hospitalization, stratified by treatment arm in HF-ACTION. Cells report mean (SD) or median [Q1,Q3].
Characteristic Usual care (n=594) Exercise training (n=561) Sample sizes and event burden (among those with $\geq 1$ recurrent event) Deaths in this subset 133 (22.4%) 122 (21.7%) Censoring (non-death) in this subset 461 (77.6%) 439 (78.3%) Hospitalizations per participant 1.66 (1.17) 1.64 (1.03) Time from last hospitalization to end-of-follow-up (days) Last hospitalization to death: median [Q1,Q3] 124 [25, 332] 165 [16, 380] Last hospitalization to death: mean (SD) 221.6 (255.7) 256.6 (289.2) Last hospitalization to censoring: median [Q1,Q3] 436 [212, 799] 494 [232, 825] Last hospitalization to censoring: mean (SD) 523.6 (373.7) 544.9 (369.7)
@@ 2 Notation, Setup, and Estimands
2.1 Potential outcomes and observed data structure
We consider a study involving $n$ units, each assigned to one of two treatment arms at the beginning of the study: $Z_{i}=1$ for the treatment group and $Z_{i}=0$ for the control group. Let $\mathbf{X}_{i}\in\mathcal{X}$ denote a vector of baseline covariates for unit $i$. For unit $i$, let $D_{i}^{z}$ denote the potential time to the terminal event under treatment $z\in\{0,1\}$. Throughout the paper, the terminal event is death; more generally, $D_{i}^{z}$ may represent any absorbing terminal event that truncates recurrent follow-up. In the motivating HF-ACTION analysis, $D_{i}^{z}$ corresponds to all-cause mortality, and the recurrent outcome consists of hospitalizations so that death serves as the truncating terminal event. Let $C_{i}^{z}$ denote the potential right-censoring time under treatment $z$ due to loss to follow-up or administrative end of study (i.e., any censoring mechanism other than death). The observed follow-up time is $\mathcal{T}_{i}=\min(D_{i}^{Z_{i}},C_{i}^{Z_{i}})$. Define the indicators $\delta_{i}^{D}=\mathbbm{1}\{\mathcal{T}_{i}=D_{i}^{Z_{i}}\}$, $\delta_{i}^{C}=\mathbbm{1}\{\mathcal{T}_{i}=C_{i}^{Z_{i}}\}$ so that $\delta_{i}^{D}=1$ indicates that death is observed and $\delta_{i}^{C}=1$ indicates right-censoring for reasons other than death.
@@ 2.1 Potential outcomes and observed data structure
We assume that the potential cumulative number of recurrent events for subject $i$ under treatment $z$, $N_{i}^{z}(\cdot)$, follows a point process. The potential point process $N_{i}^{z}(t)$ is well-defined only for $0\leq t<D_{i}^{z}$ and undefined for $t\geq D_{i}^{z}$ due to truncation; that is, we use $N_{i}^{z}(t)=*$ for the undefined case following the convention considered in Zhang and Rubin (2003). Let $N_{i}=N_{i}^{Z_{i}}(\mathcal{T}_{i}-)$ denote the observed number of recurrent events until the last follow-up. Let $T_{ij}^{z}$ represent the time of the $j$-th recurrent event for subject $i$ under treatment $z$ for any $j\in\mathbb{N}$. Since the recurrent event is truncated for those who experience the terminal event, we also define $T_{ij}^{z}=*$ for all $j$ with $T_{ij}^{z}>D_{i}^{z}$. As a special case, we write $T^{z}_{i0}=0$. The observed event times are $T_{ij}=T_{ij}^{Z_{i}}$, which are observed only for $j$ with $T_{ij}\leq\mathcal{T}_{i}$. The observed event times are $\{T_{ij}\}_{j=1}^{N_{i}}$, and the gap time between two successive events is defined as $W_{i1}^{z}=T_{i1}^{z}$ and $W_{ij}^{z}=T_{ij}^{z}-T_{i(j-1)}^{z},$ for $j\geq 2$. For the observed data, the gap times are $W_{ij}=T_{ij}-T_{i(j-1)}$. Since the time to event distribution is often right-skewed, we work on the log-time scale later in Section 3.3. Define the log-scale terminal time $U_{i}^{z}=\log(D_{i}^{z})$ and the log-scale gap times $Y_{ij}^{z}=\log(W_{ij}^{z})$ $j\in\mathbb{N}^{+}$. Because $U_{i}^{z}$ and $Y_{ij}^{z}$ are one-to-one transformations of $D_{i}^{z}$ and $W_{ij}^{z}$, modeling $(U_{i}^{z},Y_{ij}^{z})$ is equivalent to modeling $(D_{i}^{z},W_{ij}^{z})$, and follows the typical convention in the survival analysis literature. This explicitly connects the model in Section 3.3 to the causal estimands defined in Section 2.2. Finally, the observed data for each unit consist of the tuple $O_{i}=\{\mathcal{T}_{i},\delta_{i}^{D},\delta_{i}^{C},N_{i},\{T_{ij}\}_{j=1}^{N_{i}},Z_{i},\mathbf{X}_{i}\}$. We next state assumptions under which the causal estimands are identified from the observed data (up to the standard cross-world dependence addressed via sensitivity analysis in Section 3).
Assumption 1 (Stable Unit Treatment Value Assumption (SUTVA)).
There is no interference between units and no hidden versions of treatment. For each unit $i$, if $Z_{i}=z$ then $D_{i}=D_{i}^{z}$, $C_{i}=C_{i}^{z}$, and $N_{i}(t)=N_{i}^{z}(t)$ for all $t\geq 0$. Equivalently, for the transformed variables, $U_{i}=\log(D_{i})=U_{i}^{z}$ and $Y_{ij}=\log(W_{ij})=Y_{ij}^{z}$ under $Z_{i}=z$.
@@
For any follow-up time $t$, we have $Z_{i}\perp\{N_{i}^{0}(t),D_{i}^{0},N_{i}^{1}(t),D_{i}^{1}\}\mid\mathbf{X}_{i}.$
@@
Assumptions 1 and 2 are standard in randomized trials and observational studies. Assumption 1 encodes SUTVA (no interference and well-defined treatments), and Assumption 2 formulates the treatment ignorability, which is guaranteed to hold under randomization in HF-ACTION but more generally applies to observational studies with unconfounded treatment assignment. Due to the nature of the intervention and randomization, both assumptions hold in our motivating HF-ACTION trial. Assumption 3 states that potential censoring times are conditionally independent of the potential event outcomes. Although empirically unverifiable due to the cross-world counterfactuals, this assumption is reasonable in HF-ACTION because the majority of right-censored events were due to the pre-specified end of follow-up.
2.2 Principal stratum causal estimands in continuous time